OpenStudy (anonymous):

Help ~ Find the gradient of the curve at (2,1) . 2x^2 - 4xy +3y^2 = 3

OpenStudy (anonymous):

\[2x ^{2} - 4xy + 3y ^{2} = 3\]

OpenStudy (amistre64):

the gradient defines equations for the vector components of a normal to a surface

OpenStudy (amistre64):

but it has a different meaning in R^2

OpenStudy (anonymous):

ok i catch no ball .

OpenStudy (amistre64):

what is a gradient? since it has different definitions

OpenStudy (anonymous):

lols the question just wrote that. its a 4marks question of Alvl pure maths P3 !!

OpenStudy (amistre64):

calculus 1 by chance?

OpenStudy (anonymous):

answer was 2 and i dunno how the flutter it came

OpenStudy (amistre64):

please dont use that kind of language in here, its inappropriate and we have filters to catch it.

OpenStudy (amistre64):

most likely this is by implicit differentation

OpenStudy (anonymous):

ok , i apologise mate . ;D

OpenStudy (amistre64):

do you know how to take a derivative of a function?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

u meant differentiation right ? of which ones ?

OpenStudy (amistre64):

2x^2 - 4xy +3y^2 = 3 take derivatives term by term whats the derivative of 2x^2 with respect to x?

OpenStudy (anonymous):

dude please kindly take your time and write me the working very nicely. 4x

OpenStudy (amistre64):

4x is good what is the derivative of -4xy with respect to x? use the product rule

OpenStudy (anonymous):

-4y

OpenStudy (amistre64):

recall the product rule: [fg]' = f'g + fg' [-4xy]' = -4x'y -4xy' ; but x' = 1 -4y -4xy'

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

omg its getting confusing

OpenStudy (anonymous):

please just write me the working and how u got the answer please. kamsamhamnida !

OpenStudy (amistre64):

2x^2 - 4xy +3y^2 = 3; so far we got: 4x 4y -4x y' + [3y^2]' = [3]' what the derivative of 3y^2 with respect to x?

OpenStudy (anonymous):

피이팅 !

OpenStudy (amistre64):

we are alomst done, just hang in there

OpenStudy (anonymous):

hmmmm...........

OpenStudy (amistre64):

take a guess, what would this derive to if the y was an x?

OpenStudy (anonymous):

i still don't get it . =[

OpenStudy (amistre64):

what is the derivative of : 3x^2 ??

OpenStudy (anonymous):

6x

OpenStudy (amistre64):

good, but but when we dont have the x, we pop out a y' 3y^2 derives to 6y y'

OpenStudy (anonymous):

uh huh

OpenStudy (amistre64):

and the derivative of that constant is 0 as usual 2x^2 - 4xy +3y^2 = 3 derives to; 4x -4y -4x y' + 6y y' = 0 plug in your point values for x and y, and solve for y'

OpenStudy (anonymous):

oh after u get that u only sub in for x values ??

OpenStudy (amistre64):

(x=2,y=1) 4x -4y -4x y' + 6y y' = 0 4(2) -4(1) -4(2) y' + 6(1) y' = 0 8 -4 - 8y' + 6y' = 0 8 -4 + y'(-8+6) = 0 4 + y'(-2) = 0 y'(-2) = -4 y' = -4/-2 = 2

OpenStudy (amistre64):

since we have xs and ys, and the point defines our x and y; plug in and solve

OpenStudy (anonymous):

ohhh~

OpenStudy (anonymous):

so the y=2 is the final answer which is the gradient right ?

OpenStudy (amistre64):

y' = 2 is the slope of the tangent line at the given point, the gradient, yes

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

hahahs i do kind of get it , i'll try more of this questions !! THANK YOU for the HELP !! ;DD sarang hamnida.~ 사랑합니다 !

OpenStudy (amistre64):

youre welcome, and good luck ;)

OpenStudy (anonymous):

muacks!!